Two Trains leave Chicago, one traveling east at 30 miles per hour and one traveling west at 40 miles per hour. When will the trains be 280 miles apart?
Answer by stanbon!(97)

Draw the picture.
East train data: distance = x; rate=30; time = x/30
West train data: distance = 280-x; rate = 40; time = (280-x)/40
EQUATION:
Time East = Time West
x/30 = (280-x)/40
40x = 30(280-x)
40x = 30(280)-30x
70x =30(280
x = 3(40)
x = 120 mph
Therefore time = x/30 = 120/30= 4 hours
Cheers,
Stan H.